# The discovery and characterization of minimoon 2024 PT5

### v1

@@ Abstract
Minimoons are asteroids that become temporarily captured by the Earth-Moon system. We present the discovery of 2024 PT5, a minimoon discovered by the Asteroid Terrestrial-impact Last Alert System (ATLAS) Sutherland telescope on 2024 August 7. The minimoon with heliocentric semi-major axis, $a$$\sim$1.01 au, and perihelion, $q$$\sim$0.99 au, became captured by the Earth-Moon system on 2024 September 29 and left on 2024 November 25 UTC. Visible g, r, i, and Z spectrophotometry was obtained using Gemini North/Gemini Multi-Object Spectrograph (GMOS) on 2024 September 27. The color indices are g-r = 0.58$\pm$0.04, r-i = 0.29$\pm$0.04, i-Z = -0.27$\pm$0.06, and the spectrum best matches lunar rock samples followed by S-complex asteroids. Assuming an albedo of 0.21 and using our measured absolute magnitude of 28.64$\pm$0.04, 2024 PT5 has a diameter of 5.4$\pm$1.3 m. We also detect variations in the lightcurve of 2024 PT5 with a 0.28$\pm$0.07 magnitude amplitude and a double-peaked period of $\sim$2600$\pm$500 s. We improve the orbital solution of 2024 PT5 with our astrometry and estimate the effect of radiation pressure on its deriving an area-to-mass ratio of 7.02$\pm$2.05$\times$10-5 m2/kg, implying a density of $\sim$3.9$\pm$2.1 g/cm3, compatible with having a rocky composition. If we assume 2024 PT5 is from the NEO population, its most likely sources are resonances in the inner Main Belt by comparing its orbit with the NEO population model, though this does not exclude a lunar origin.
@@ 1 Introduction
Evidence from observations of near-Earth objects (NEOs) suggests that the majority originate from the Main Belt asteroids (MBAs) and Jupiter Family Comet (JFC) populations (Granvik et al., 2018; Nesvorný et al., 2023, 2024). A small percentage of NEOs can become temporary Earth co-orbitals, located in a 1:1 mean motion resonance with the Earth (Morais & Morbidelli, 2002) where they can become Trojans or quasi-satellites of the Earth for several thousand years (Brasser et al., 2004). A small portion of NEOs, less than 1/10,000,000 of the total NEO population, can become temporarily captured by the Earth-Moon system’s gravity and enter its Hill sphere, becoming a so-called “minimoon” (Granvik et al., 2013; Jedicke et al., 2018).
@@ 2 Observations
The discovery observations of 2024 PT5 were taken on 2024 August 7 21:11:57 UTC at the ATLAS Sutherland telescope (Minor Planet Center (MPC) observing code M22, Fig. 1, Williams, 2024). At the time of the asteroid’s discovery, the asteroid was located at right ascension (RA) = 18 57 04.97, declination = -55 56 00.24 and 1.54 lunar distances (0.00397 au) au from the Earth, a heliocentric distance of 1.017 au, at a phase angle of 128.747, and had a visible magnitude of $\sim$17. The asteroid was detected in four 30 s ATLAS o-band filter exposures spaced over $\sim$1700 s. The ATLAS o-band filter provides coverage between 560 nm and 820 nm with an effective wavelength of 663 nm (Tonry et al., 2018). The asteroid was moving 13.7/d (0.57″/s) in the northwest direction at the time of discovery, resulting in having a trailed appearance in the ATLAS images (panels a, b, c, and d of Fig 1). The trailed exposures of 2024 PT5 show little to no curvature over the 1700 s span of the discovery detections with a great circle residual of $\sim$1.2″.
@@
On 2024 September 27 05:54 UTC, 2024 PT5 was observed using the 8.1 m Gemini North telescope with the Gemini Multi-Object Spectrograph (GMOS) instrument (Hook et al., 2004) in imaging mode (Program ID GN-2024B-FT-106, PI B. Bolin). GMOS was used with the Hamamatsu array with an effective pixel scale of 0.0807″/pixel in $2\times 2$ binning mode. The images were taken in Sloan Digital Sky Survey (SDSS)-equivalent g, r, and i filters (effective wavelengths 0.475 microns, 0.630 microns and 0.780 microns, Fukugita et al., 1996). For coverage of the spectrum of 2024 PT5 past 0.8 microns, we use the GMOS-N Z-band filter, which is effectively equivalent to the United Kingdom Infrared Telescope Wide-Field Camera (WFCAM) Z filter (Hewett et al., 2006). The WFCAM Z filter (effective wavelength 0.880 microns, Casali et al., 2007) was designed to be equivalent to the SDSS z filter (effective wavelength 0.89 microns) but avoids the atmospheric absorption band near 0.95 microns (Hodgkin et al., 2009).
@@ 3.1 Astrometry and orbital determination
The astrometry measured from the GMOS images was submitted to the MPC (Denneau et al., 2024). The GMOS astrometry of 2024 PT5 was combined with 200 additional 2024 PT5 observations taken independently by follow-up observers between 2024 August 7 UTC and 2024 October 24 UTC from the MPC archive [^2] https://minorplanetcenter.net/db_search/show_object?utf8=%E2%9C%93&object_id=2024+PT5, accessed on 2024 November 7.. A complete table of the astrometry of 2024 PT5 is provided in Table A1 of the appendix. We adopt conservative estimates for the astrometric uncertainties of $\sim$1.0″ in both right ascension and declination for the other observatories’ measured positions except for observations made at the Great Shefford Observatory in West Berkshire, England (MPC observatory code J95), where we used an estimated uncertainty of 0.4″ (Vereš et al., 2017).
Using the Find__\__Orb orbital determination software by Bill Gray [^3] https://www.projectpluto.com/find_orb.htm, we fit an orbit to the observations in Table A1, using the 8 planets and the Moon as perturbers, and including the effects of solar radiation pressure (e.g., Mommert et al., 2014; Fedorets et al., 2020). Our orbital fit results in estimates of the heliocentric and geocentric semi-major axis, $a$, eccentricity, $e$, inclination, $i$, ascending node, $\Omega$, argument of perihelion, $\omega$, mean anomaly, $M$, and the area-to-mass ratio (AMR) as a measure of the effect of solar radiation pressure on the orbit of 2024 PT5.
@@ 3.2 Photometry and spectral classification
Figure 4: a vs. i-z(Z) colors of 2024 PT5 plotted with a vs. i-z colors of C, S and V type asteroids from (Ivezić et al., 2001; Jurić et al., 2002), active comets (Solontoi et al., 2012), and Kuiper Belt Objects (Ofek, 2012). The colorization scheme of data points as a function of a and i-z is adapted from Ivezić et al. (2002) where blue symbol colors correspond to C-type asteroids, red symbol colors correspond to S-type asteroids and green symbol colors correspond to V-type asteroids. We note that in this case the measured Z magnitude of 2024 PT5 is plotted as a substitute for its z magnitude. $i$-$z$. The a and $i$-$z$ range of average S, V, and C-type asteroids are shown computed from the average spectra from DeMeo et al. (2009).
We compute the spectral reflectance of 2024 PT5 by dividing its flux per g, r, i, and Z filter obtained by the flux of a solar analog in a corresponding filter, normalizing it to a wavelength of 550 nm. The normalized reflectivity spectrum of 2024 PT5 shown in Fig. 5 is similar to the spectra of S- and V-type asteroids (Bus & Binzel, 2002; DeMeo et al., 2009) with a red slope between 0.47 microns and 0.77 microns and a decrease in reflectance near 0.88 microns. The spectrum of 2024 PT5 is also similar to some bulk basaltic lunar rock samples consisting of pyroxene minerals (Isaacson et al., 2011). We compute the $\chi^{2}$ statistic for the spectrum of 2024 PT5 compared to the spectrum of 19 different asteroid types including S-complex asteroid types (A, S, Sq, Sv, O, Q, R, V), C-complex types (B, C, Cg, Cgh, D), and X-complex types (K, L, X, Xc, Xe, Xk, Xn) from (DeMeo et al., 2009). We find the closest match is to lunar rock with a reduced $\chi^{2}$ of 2.76 and Sv-type asteroids with a reduced $\chi^{2}$ of 3.75.
@@ 3.3 Lightcurve, periodicity, and axial ratio
The Lomb-Scargle periodogram (Lomb, 1976) was applied to the time series g filter data, as seen in the top panel of Fig. 7. The highest peak in the lightcurve period vs. spectral power curve is located at $\sim$1320 s with a formal significance of $p\simeq$ 3 x 10-4. We apply bootstrap estimation (Press et al., 1986) of the uncertainties by removing $\sqrt{N}$ data points from the time series lightcurve and repeating our periodogram estimation of the lightcurve period 10,000 times, resulting in a central value of $\sim$1317 s and a 1 $\sigma$ uncertainty estimate of $\sim$227 s. We apply phase dispersion minimization analysis to our data (Stellingwerf, 1978) as an independant check of the Lomb-Scargle results, obtaining a result of $\sim$1150 s as seen in the bottom panel of Fig. 7, comparable with the lightcurve period estimate. These period estimates imply that 2024 PT5 has a double-peaked rotation period of $\sim$2600$\pm$500 s.
@@ 4 Discussion and Conclusions
We use the albedo estimate and H magnitude of 2024 PT5 to determine the diameter, D, of 2024 PT5 using the equation $\mathrm{D=\frac{1329}{\sqrt{p_{v}}}10^{-\frac{H}{5}}}$ from Russell (1916) arriving at D = 5.4$\pm$1.2 m. Combining this diameter calculation with the AMR determined from our orbital fit, we can estimate a bulk density for 2024 PT5 of 3.9$\pm$2.1g/cm3 comparable with achondritic basaltic meteorites (Macke et al., 2011) and small asteroids with AMR measurements (Mommert et al., 2014). If we use the mean albedo of lunar surface samples, 0.14 (Matthews, 2008), we obtain an diameter of 6.0$\pm$1.7 m and a bulk density of 3.6$\pm$2.4g/cm3. This density estimate is also comparable to the density of some lunar rock samples (Kiefer et al., 2012). Using this these density estimates, the mass of 2024 PT5 is $\sim$105 kg. With the comparison of its physical properties determined from our observations to the physical properties of asteroids and lunar samples, we conclude that 2024 PT5 the physical properties of 2024 PT5 deduced from our observations is compatible within an inner Main Belt or lunar ejecta origin.

### v2

@@ Abstract
Minimoons are asteroids that become temporarily captured by the Earth-Moon system. We present the discovery of 2024 PT5, a minimoon discovered by the Asteroid Terrestrial-impact Last Alert System (ATLAS) Sutherland telescope on 2024 August 7. The minimoon with heliocentric semi-major axis, $a$$\sim$1.01 au, and perihelion, $q$$\sim$0.99 au, became captured by the Earth-Moon system on 2024 September 29 and left on 2024 November 25 UTC. Visible g, r, i, and Z spectrophotometry was obtained using Gemini North/Gemini Multi-Object Spectrograph (GMOS) on 2024 September 27. The color indices are g-r = 0.58$\pm$0.04, r-i = 0.29$\pm$0.04, i-Z = -0.27$\pm$0.06, and the spectrum best matches lunar rock samples followed by S-complex asteroids. Assuming an albedo of 0.21 and using our measured absolute magnitude of 28.64$\pm$0.04, 2024 PT5 has a diameter of 5.4$\pm$1.2 m. We also detect variations in the lightcurve of 2024 PT5 with a 0.28$\pm$0.07 magnitude amplitude and a double-peaked period of $\sim$2600$\pm$500 s. We improve the orbital solution of 2024 PT5 with our astrometry and estimate the effect of radiation pressure on its deriving an area-to-mass ratio of 7.02$\pm$2.05$\times$10-5 m2/kg, implying a density of $\sim$3.9$\pm$2.1 g/cm3, compatible with having a rocky composition. If we assume 2024 PT5 is from the NEO population, its most likely sources are resonances in the inner Main Belt by comparing its orbit with the NEO population model, though this does not exclude a lunar origin.
@@ 1 Introduction
Evidence from observations of near-Earth objects (NEOs) suggests that the majority originate from the Main Belt asteroids (MBAs) and Jupiter Family Comet (JFC) populations (Granvik et al., 2018; Nesvorný et al., 2023, 2024). A small percentage of NEOs can become temporary Earth co-orbitals, located in a 1:1 mean motion resonance with the Earth (Morais & Morbidelli, 2002) where they can become Trojans or quasi-satellites of the Earth for several thousand years (Brasser et al., 2004). A small portion of NEOs, less than $<10^{-7}$ of the total NEO population, can become temporarily captured by the Earth-Moon system’s gravity and enter its Hill sphere, becoming a so-called “minimoon” (Granvik et al., 2013; Jedicke et al., 2018).
@@ 2 Observations
The discovery observations of 2024 PT5 were taken on 2024 August 7 21:11:57 UTC at the ATLAS Sutherland telescope, Minor Planet Center (MPC) observing code M22, (Fig. 1, Denneau et al., 2024). At the time of the asteroid’s discovery, the asteroid was located at right ascension (RA) = 18 57 04.97, dec = -55 56 00.24 and 1.54 lunar distances (0.00397 au) au from the Earth, a heliocentric distance of 1.017 au, at a phase angle of 128.747, and had a visible magnitude of $\sim$17. The asteroid was detected in four 30 s ATLAS o-band filter exposures spaced over $\sim$1700 s. The ATLAS o-band filter provides coverage between 560 nm and 820 nm with an effective wavelength of 663 nm (Tonry et al., 2018). The asteroid was moving 13.7/d (0.57″/s) in the northwest direction at the time of discovery, resulting in having a trailed appearance in the ATLAS images (panels a, b, c, and d of Fig 1). The trailed exposures of 2024 PT5 show little to no curvature over the 1700 s span of the discovery detections with a great circle residual of $\sim$1.2″.
@@
On 2024 September 27 05:54 UTC, 2024 PT5 was observed using the 8.1 m Gemini North telescope with the Gemini Multi-Object Spectrograph (GMOS) instrument (Hook et al., 2004) in imaging mode (Program ID GN-2024B-FT-106, PI B. Bolin). GMOS was used with the Hamamatsu array with an effective pixel scale of 0.0807″/pixel in $2\times 2$ binning mode. The images were taken in Sloan Digital Sky Survey (SDSS)-equivalent g, r, and i filters (effective wavelengths 475 nm, 630 nm, and 780 nm, Fukugita et al., 1996). For coverage of the spectrum of 2024 PT5 past 800 nm, we use the GMOS-N Z-band filter, which is effectively equivalent to the United Kingdom Infrared Telescope Wide-Field Camera (WFCAM) Z filter (Hewett et al., 2006). The WFCAM Z filter (effective wavelength 880 nm, Casali et al., 2007) was designed to be equivalent to the SDSS z filter (effective wavelength 890 nm) but avoids the atmospheric absorption band near 950 nm (Hodgkin et al., 2009).
@@ 3.1 Astrometry and orbital determination
The astrometry measured from the GMOS images was submitted to the MPC (Williams, 2024). The GMOS astrometry of 2024 PT5 was combined with 200 additional 2024 PT5 observations taken independently by follow-up observers between 2024 August 7 UTC and 2024 October 24 UTC from the MPC archive [^2] https://minorplanetcenter.net/db_search/show_object?utf8=%E2%9C%93&object_id=2024+PT5, accessed on 2024 November 7.. A complete table of the astrometry of 2024 PT5 is provided in Table A1 of the appendix. We adopt conservative estimates for the astrometric uncertainties of $\sim$1.0″ in both right ascension and declination for the other observatories’ measured positions except for observations made at the Great Shefford Observatory in West Berkshire, England (MPC observatory code J95), where we used an estimated uncertainty of 0.4″ (Vereš et al., 2017).
Using the Find__\__Orb orbital determination software by Bill Gray [^3] https://www.projectpluto.com/find_orb.htm, we fit an orbit to the observations in Table A1, using the 8 planets and the Moon as perturbers, and including the effects of solar radiation pressure (e.g., Fedorets et al., 2020; Jewitt, 2024). Our orbital fit results in estimates of the heliocentric and geocentric semi-major axis, $a$, eccentricity, $e$, inclination, $i$, ascending node, $\Omega$, argument of perihelion, $\omega$, mean anomaly, $M$, and the area-to-mass ratio (AMR) as a measure of the effect of solar radiation pressure on the orbit of 2024 PT5.
@@ 3.2 Photometry and spectral classification
Figure 4: a vs. i-z(Z) colors of 2024 PT5 plotted with a vs. i-z colors of C, S and V type asteroids from (Ivezić et al., 2001; Jurić et al., 2002), active comets (Solontoi et al., 2012), Kuiper Belt Objects (Ofek, 2012), and bulk lunar rock samples (Isaacson et al., 2011). The colorization scheme of data points as a function of a and i-z is adapted from Ivezić et al. (2002) where blue symbol colors correspond to C-type asteroids, red symbol colors correspond to S-type asteroids and green symbol colors correspond to V-type asteroids. We note that in this case the measured Z magnitude of 2024 PT5 is plotted as a substitute for its z magnitude. $i$-$z$. The a and $i$-$z$ range of average S, V, and C-type asteroids are shown computed from the average spectra from DeMeo et al. (2009).
We compute the spectral reflectance of 2024 PT5 by dividing its flux per g, r, i, and Z filter obtained by the flux of a solar analog in a corresponding filter, normalizing it to a wavelength of 550 nm. The normalized reflectivity spectrum of 2024 PT5 shown in Fig. 5 is similar to the spectra of S- and V-type asteroids (Bus & Binzel, 2002; DeMeo et al., 2009) with a red slope between 470 nm and 770 nm and a decrease in reflectance near 880 nm. The spectrum of 2024 PT5 is also similar to some bulk basaltic lunar rock samples consisting of pyroxene minerals (Isaacson et al., 2011). For a thorough comparison, we compute the $\chi^{2}$ statistic for the spectrum of 2024 PT5 compared to the spectrum of 19 different asteroid types including S-complex asteroid types (S, Sq, Sv, Q), C-complex types (B, C, Cg, Cgh), X-complex types (X, Xc, Xe, Xk, Xn), and other assorted types (A, D, K, L, O, R, V) from (Bus & Binzel, 2002; DeMeo et al., 2009). We find the closest match is to lunar rock with a reduced $\chi^{2}$ of 2.76 and Sv-type asteroids with a reduced $\chi^{2}$ of 3.75. By comparison, the comparison with V-type spectra results in a reduced $\chi^{2}$ of $\sim$12.84.
@@ 3.3 Lightcurve, periodicity, and axial ratio
The Lomb-Scargle periodogram (Lomb, 1976) was applied to the time series g filter data, as seen in the top panel of Fig. 7. We take a similar approach as Bolin et al. (2024) to estimating the uncertainty on the determined period by removing $\sqrt{N}$ data points from the time series lightcurve and repeating our periodogram estimation of the lightcurve period 10,000 times, resulting in a central value of $\sim$1317 s and a 1 $\sigma$ uncertainty estimate of $\sim$227 s. We apply phase dispersion minimization analysis to our data (Stellingwerf, 1978) as an independant check of the Lomb-Scargle results, obtaining a result of $\sim$1150 s as seen in the bottom panel of Fig. 7, comparable with the lightcurve period estimate. These period estimates imply that 2024 PT5 has a double-peaked rotation period of $\sim$2600$\pm$500 s.
@@ 4 Discussion and Conclusions
We use the albedo estimate and H magnitude of 2024 PT5 to determine the diameter, D, of 2024 PT5 using the equation $\mathrm{D=\frac{1329}{\sqrt{p_{v}}}10^{-\frac{H}{5}}}$ from Russell (1916) arriving at D = 5.4$\pm$1.2 m. Combining this diameter calculation with the AMR determined from our orbital fit, we can estimate a bulk density for 2024 PT5 of 3.9$\pm$2.1g/cm3 comparable with achondritic basaltic meteorites (Macke et al., 2011) and small asteroids with AMR measurements (Mommert et al., 2014). If we use the mean albedo of lunar rock samples of 0.14, we obtain an diameter of 6.0$\pm$1.7 m and a bulk density of 3.6$\pm$2.4g/cm3. This density estimate is comparable to the density of some lunar rock samples (Kiefer et al., 2012). Using these density estimates, the mass of 2024 PT5 is $\sim$105 kg. With the comparison of its physical properties determined from our observations to the physical properties of asteroids and lunar samples, we conclude that 2024 PT5 the physical properties of 2024 PT5 deduced from our observations is compatible within an inner Main Belt or lunar ejecta origin.